OpenStudy (anonymous):

(5/x)+2=(5/3x)+(13/6) What are the restrictions on the variable? What is the answer to the equation?

OpenStudy (anonymous):

x cannot equal 0

OpenStudy (anonymous):

restriction: x not equal 0, solution: 5/x-5/3x=13/6-2/1 15/3x-5/3x=13/6-12/6 10/3x=1/6 10*6=3x*1 60=3x x=20 Check: 5/20+2=5/60+13/6 5/20+40/20=5/60+130/60 45/20=135/60 true